# SMURF: Continuous Dynamics for Motion-Deblurring Radiance Fields

### v1

@@ 4.1 Implementation Details
In this experiment, we utilize the camera motion blur dataset published by [31], which is divided into synthetic and real-world scenes. The synthetic scene dataset comprises five scenes synthesized using Blender [8], with multi-view cameras set up to render images with applied camera motion. The camera motion trajectory is formed by linearly interpolating between the poses of the first and last cameras, and the rendered multi-view images are combined in the linear RGB space to obtain the final blurry images. The real-world scene dataset consists of ten scenes captured with an actual camera. The blurry images are obtained by physically shaking the camera during the exposure time, and the camera poses are calculated using the images obtained with COLMAP [46, 45].
Methods Synthetic Scene Dataset Real-World Scene Dataset  Training Time (h)  Rendering Time (s)
Naive NeRF [33] 23.78 0.6807 0.3362 22.69 0.6347 0.3687 $\sim$ 6 $\sim$ 7
NeRF+MPR [60] 25.11 0.7476 0.2148 23.38 0.6655 0.3140 $\sim$ 6 $\sim$ 7
NeRF+PVD [47] 24.58 0.7190 0.2475 23.10 0.6389 0.3425 $\sim$ 6 $\sim$ 7
Deblur-NeRF [31] 28.77 0.8593 0.1400 25.63 0.7645 0.1820 $\sim$ 25 $\sim$ 7
PDRF-10* [40] 28.86 0.8795 0.1139 25.90 0.7734 0.1825 2.9 0.91
BAD-NeRF* [56] 27.32 0.8178 0.1127 22.82 0.6315 0.2887 $\sim$ 20 $\sim$ 7
DP-NeRF [25] 29.23 0.8674 0.1184 25.91 0.7751 0.1602 $\sim$ 26 $\sim$ 7
SMURF 30.98 0.9147 0.0609 26.52 0.7986 0.1013 1.7 0.77
Table 1: Quantitative comparisons on synthetic and real-world scene dataset. We evaluate the quantitative performance using three metrics, demonstrating that our proposed model significantly ourperforms existing ones even with faster training and rendering times. The red and orange cells respectively indicate the highest and second-highest value.
Figure 4: Qualitative comparisons on synthetic scenes and real-world scenes. Without long training times and sharp images, SMURF produces results most similar to the reference images. SMURF models the detailed aspects that are not captured by previous methods with long training time.
Training Details.
We implement our model on TensoRF [6], an explicit grid-based method, as our backbone renderer. We upsample the voxel counts of grids $\mathcal{G}_{\sigma}$ and $\mathcal{G}_{c}$ from $128^{3}$ to $480^{3}$, and set the feature component $F$ of the grids to 36 and 96, respectively. To implement the CMBK, we set the number of warped rays $\mathcal{N}$ on the camera motion trajectory to 8 and adopt the Euler method [12] as the solver for the Neural-ODE. Single scene training is conducted for 40k iterations, with the learning rates for the grids decaying from $0.02$ to $0.002$, and the learning rate for CMBK decaying from $0.001$ to $0.0001$. The weights for the losses $\lambda_{TV}^{\sigma}$, $\lambda_{TV}^{c}$, and $\lambda_{supp}$ are all set to 0.1. All our experiments are conducted on a single NVIDIA RTX 3090.
Evaluation Metrics.
The images rendered through the SMURF are quantitatively evaluated using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS).
Synthetic Scene Factory CozyRoom Pool Tanabata Trolley
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
Naive NeRF 19.32 0.4563 0.5304 25.66 0.7941 0.2288 30.45 0.8354 0.1932 22.22 0.6807 0.3653 21.25 0.6370 0.3633
MPR+NeRF 21.70 0.6153 0.3094 27.88 0.8502 0.1153 30.64 0.8385 0.1641 22.71 0.7199 0.2509 22.64 0.7141 0.2344
PVD+NeRF 20.33 0.5386 0.3667 27.74 0.8296 0.1451 27.56 0.7626 0.2148 23.44 0.7293 0.2542 23.81 0.7351 0.2567
Deblur-NeRF 25.60 0.7750 0.2687 32.08 0.9261 0.0477 31.61 0.8682 0.1246 27.11 0.8640 0.1228 27.45 0.8632 0.1363
PDRF-10* 25.87 0.8316 0.1915 31.13 0.9225 0.0439 31.00 0.8583 0.1408 28.01 0.8931 0.1004 28.29 0.8921 0.0931
BAD-NeRF* 24.43 0.7274 0.2134 29.77 0.8864 0.0616 31.51 0.8620 0.0802 25.32 0.8081 0.1077 25.58 0.8049 0.1008
DP-NeRF 25.91 0.7787 0.2494 32.65 0.9317 0.0355 31.96 0.8768 0.0908 27.61 0.8748 0.1033 28.03 0.8752 0.1129
SMURF 29.87 0.8958 0.1057 32.48 0.9285 0.0379 32.34 0.8884 0.0779 29.91 0.9300 0.0436 30.30 0.9307 0.0397
Table 2: Comparison of performance for individual scenes on the synthetic dataset. SMURF exhibits higher performance across all scenes, with the exception of “CozyRoom,” where it shows slightly lower performance relative to others.
Real-World Scene Ball Basket Buick Coffee Decoration
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
Naive NeRF 24.08 0.6237 0.3992 23.72 0.7086 0.3223 21.59 0.6325 0.3502 26.48 0.8064 0.2896 22.39 0.6609 0.3633
Deblur-NeRF 27.36 0.7656 0.2230 27.67 0.8449 0.1481 24.77 0.7700 0.1752 30.93 0.8981 0.1244 24.19 0.7707 0.1862
PDRF-10* 27.37 0.7642 0.2093 28.36 0.8736 0.1179 25.73 0.7916 0.1582 31.79 0.9002 0.1133 23.55 0.7508 0.2145
BAD-NeRF* 21.33 0.5096 0.4692 26.44 0.8080 0.1325 21.63 0.6429 0.2593 28.98 0.8369 0.1956 22.13 0.6316 0.2894
DP-NeRF 27.20 0.7652 0.2088 27.74 0.8455 0.1294 25.70 0.7922 0.1405 31.19 0.9049 0.1002 24.31 0.7811 0.1639
SMURF 27.50 0.7760 0.1298 28.95 0.8842 0.0619 27.10 0.8409 0.0839 31.33 0.8879 0.0874 24.90 0.8114 0.1190
Real-World Scene Girl Heron Parterre Puppet Stair
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
Naive NeRF 20.07 0.7075 0.3196 20.50 0.5217 0.4129 23.14 0.6201 0.4046 22.09 0.6093 0.3389 22.87 0.4561 0.4868
Deblur-NeRF 22.27 0.7976 0.1687 22.63 0.6874 0.2099 25.82 0.7597 0.2161 25.24 0.7510 0.1577 25.39 0.6296 0.2102
PDRF-10* 24.12 0.8328 0.1679 22.53 0.6880 0.2358 25.36 0.7601 0.2263 25.02 0.7496 0.1532 25.20 0.6235 0.2288
BAD-NeRF* 18.10 0.5652 0.3933 22.18 0.6479 0.2226 23.44 0.6243 0.3151 22.48 0.6249 0.2762 21.52 0.4237 0.3341
DP-NeRF 23.33 0.8139 0.1498 22.88 0.6930 0.1914 25.86 0.7665 0.1900 25.25 0.7536 0.1505 25.59 0.6349 0.1772
SMURF 25.66 0.8592 0.0829 23.59 0.7317 0.1381 25.47 0.7825 0.1207 25.19 0.7702 0.1077 25.48 0.6421 0.0822
Table 3: Comparison of performance for individual scenes on the real-world dataset. While SMURF shows slightly lower PSNR and SSIM in some scenes, it shows significantly higher LPIPS across all scenes compared to previous methods.
@@ 4.2 Novel View Synthesis Results
Quantitative Results.
We show the quantitative evaluation of our network, SMURF, in comparison to various baselines on the two benchmark datasets proposed by Deblur-NeRF [31], as shown in Tab. 1. NeRF+MPR and NeRF+PVD are trained with a naive NeRF, where the single-image deblurring methods MPR [60] and PVD [47] to the input data. Additionally, since PDRF-10 [40] does not specify LPIPS in their experiment, we obtain performance of three evaluation metrics using their released code. Furthermore, as BAD-NeRF [56] specifies results from experiments on a modified benchmark dataset, for a fair evaluation, we apply the released code of BAD-NeRF to the benchmark dataset. According to  Tab. 1, despite significantly reduced training and rendering times compared to previous methods, we demonstrate superior quantitative performance across all the metrics. Especially, LPIPS of SMURF demonstrates SMURF’s exceptional perceptual quality across all datasets. Tab. 2 and Tab. 3 detail the individual performance of scenes within the synthetic and real-world datasets, respectively. For synthetic scenes, except for “Cozyroom,” all scenes show quantitatively high performance, with this scene also displaying no significant difference when compared to DP-NeRF. Additionally, for the real-world scenes, despite a few scenes exhibiting somewhat lower PSNR, they demonstrate better perceptual quality through superior LPIPS scores.
Qualitative Results.
We validate the effectiveness of SMURF’s quantitatively high performance through qualitative evaluation via novel view rendering. According to Fig. 4, we compare our results with previous 3D scene deblurring methods for two synthetic scenes (“Factory” and “Tanabata”) and three real-world scenes (“Girl”, “Buick”, and “Basket”). For the “Factory” scene, when comparing the perceptual quality of the reconstructed lowest part of the stairs, it is evident that BAD-NeRF [56] and SMURF best estimate the 3D geometry. However, BAD-NeRF shows an inability to capture the overall color as accurately as other methods. In the rendering results for the “Tanabata” scene, while other methods fail to accurately estimate the position of the power lines, SMURF restores them most similarly to the reference image. In the “Girl” and “Buick” scenes, our method notably restores the visual quality of the shelf’s black support rods and the car logo, respectively, most closely to the reference images. For the “Basket” scene, the holes of the basket further demonstrate the superior qualitative outcome of our results. For additional novel view rendering results, refer to our Appendix.
@@ 4.3 Ablation Study
To demonstrate the effectiveness of the chrono-view embedding function $\mathbf{\Psi}$ in making continuous latent space modeling robust to exposure time, as discussed in Sec. 3, we conduct various ablative experiments. We conduct experiments by dividing the embedding types into three categories. The time embedding assumes that all images have the same exposure time without incorporating view information, while the chrono-time embedding applies view information, ensuring that all images have distinct exposure time information. As shown in Tab. 4, the application of time embedding yields higher performance in general, but chrono-time embedding shows higher performance only in real-world scenes, while its effect in synthetic scenes is minimal. This is attributed to the characteristics of the synthetic dataset created by Deblur-NeRF [31] using Blender [8], where motion blur images are acquired by linearly interpolating between camera poses to synthesize a total of 10 images. In other words, the synthetic scenes set a constant speed for camera motion throughout the exposure time for all images, implying that all input images share the same exposure time information. Our chrono-view embedding function assumes that all images have varying exposure times, leading to its minimal effect on the synthetic dataset. However, the real-world dataset consists of images captured with an actual camera, where exposure times are likely to vary unless shutter speed is fixed, and the speed of camera motion is not constant during the exposure time. Therefore, applying this function to the real-world scenes yield improved performance due to these variations.
Methods  Embedding Synthetic Scene Dataset Real-World Scene Dataset
Type  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)
SMURF None 30.54 0.8995 0.0709 25.74 0.7755 0.1194
SMURF Time 30.94 0.9158 0.0610 26.34 0.7881 0.1102
SMURF Chrono-View 30.98 0.9147 0.0609 26.52 0.7986 0.1013
Table 4: Ablations on chrono-view embedding function. We conduct an ablation study by dividing the embedding types into three categories: no embedding, time embedding, and chrono-time embedding.
@@
To validate the effectiveness of the regularization strategies presented in Sec. 3, namely output suppression loss and residual momentum, we conduct various experiments as shown in Tab. 5. The output suppression loss is designed to suppress changes in the origin and direction of the initial ray on the camera motion trajectory to zero, yielding higher performance when applied. This is because the changes in rays are prevented from diverging since the estimated poses also share the same derivative function $f$, similar to the effect of the alignment loss proposed by Deblur-NeRF [31]. This similarity suggests that aligning ray origins is robust to inaccuracies in camera pose, which can be ambiguous due to motion blur. Consequently, Tab. 5 demonstrates that the performance of PSNR and SSIM in synthetic and real-world scenes improves more significantly.
Methods  O.S. Loss Residual Synthetic Scene Dataset Real-World Scene Dataset
 Momentum  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)
SMURF 26.17 0.8186 0.0759 25.27 0.7576 0.1121
SMURF 28.49 0.8727 0.0675 25.60 0.7748 0.1057
SMURF 30.72 0.9052 0.0684 25.99 0.7846 0.1181
SMURF 30.98 0.9147 0.0609 26.52 0.7986 0.1013
Table 5: Ablations on regularization strategies. We conduct ablative experiments on the two proposed normalization methods. “O.S. Loss” refers to the output suppression loss.
Residual momentum offers a similar effect to output suppression loss. While the change in camera poses is estimated through a parameterized differential equation, the shape of the derivative obtained in an unsupervised manner could diverge. Therefore, we use residual momentum to ensure the next pose does not deviate significantly from the previous one. Tab. 5 shows that applying another regularization strategy, residual momentum, shows better performance in terms of LPIPS than output suppression loss. Combining two strategies, the SMURF results in best performance across all metrics.
@@ 5 Conclusion
We have proposed SMURF, a novel approach that sequentially models accurate camera motion for reconstructing sharp 3D scenes from motion-blurred images. Unlike previous approaches that estimate camera motion in a single step, SMURF incorporates, for the first time, a kernel for estimating sequential camera motions, named the CMBK. This camera motion is represented with continuity by solving continuous dynamics in the latent space using Neural-ODEs. To prevent the divergence of rays estimated by CMBK beyond the motion blur range, we apply regularization techniques: residual momentum and output suppression loss. Furthermore, we model the 3D scene using tensor factorization-based representation, which allows for the integration of incomplete blurred information and complete sharp information within adjacent voxels via CMBK, thereby reducing the uncertainty of blurry information. SMURF significantly outperform previous works quantitatively with faster training and rendering, and its qualitative evaluation is demonstrated through novel view rendering results.
Appendix
Appendix A Details of CMBK
Figure 5: Detailed process for generating kernel. To highlight the continuous dynamics in the latent space, all embedding functions are omitted from the figure. $\mathbf{p}(x_{t_{0}})$ and $\mathbf{p}(y_{t_{0}})$ correspond to the $x$ and $y$ coordinates of the pixel associated with the initial ray, respectively.
We elaborate on the continuous dynamics of the proposed CMBK as shown in Fig. 5. We assume that camera motion encompasses inherent dynamics with a unique solution. The assumed solution is represented by the lighter circles in the left plots of Fig. 5 (b). Rather than implementing the inherent dynamics in a simple physical space, we refine them within a latent space with parametric learning. The continuous dynamics of CMBK involve transforming the pixel coordinates corresponding to the ray into latent features via a parameterized encoder $\mathcal{E}_{\theta}$, and a unique numerical solution [29] is obtained by solving the initial value problem on the formed latent space through a neural ordinary differential equation [7]. The solution obtained represents the change in pixel of the warped ray relative to the initial ray’s pixel. These changes define the pixels corresponding to the warped rays, and we specify Eq. (11) from the main paper:
$\begin{array}[]{ccc}\vspace{+1mm}\mathbf{p}^{t_{i}}=\left(\mathbf{p}(x_{t_{i}}),~{}\mathbf{p}(y_{t_{i}})\right),\\ \vspace{+1mm}\mathbf{p}(x_{t_{i+1}})=\mathbf{p}(x_{t_{i}})+\Delta\mathbf{p}(x_{t_{i+1}}),\\ \mathbf{p}(y_{t_{i+1}})=\mathbf{p}(y_{t_{i}})+\Delta\mathbf{p}(y_{t_{i+1}}).\end{array}$ (15)
Following this principle, the change in the initial ray must necessarily be zero, and to ensure this, the proposed method is the output suppression loss.
Appendix B Number of Warped Rays
Methods  $\mathcal{N}$ Synthetic Scene Dataset Real-World Scene Dataset  Training Time (h)
SMURF 4 29.52 0.8829 0.1014 25.81 0.7710 0.1388 1.16
SMURF 5 30.30 0.9013 0.0926 25.91 0.7811 0.1253 1.27
SMURF 6 29.85 0.8929 0.0851 25.98 0.7822 0.1155 1.43
SMURF 7 30.23 0.8964 0.0751 26.40 0.7944 0.1052 1.56
SMURF 8 30.98 0.9147 0.0609 26.52 0.7986 0.1013 1.72
SMURF 9 30.41 0.9086 0.0575 26.24 0.7922 0.1021 1.88
Table 6: Performance and training time of SMURF according to the number of warped rays. The red, orange, and yellow cells respectively indicate the highest, second-highest, and third-highest value.
We conduct extensive experiments to analyze the performance of the proposed CMBK based on the number of warped rays, $\mathcal{N}$. As shown in Tab. 6, larger $\mathcal{N}$ requires the more pixels to be rendered, resulting in an almost linear increase in training time. Moreover, Fig. 6 shows the performance according to the number of warped rays. Across all datasets, an increase in the number of warped rays tends to improve the PSNR and SSIM metrics. LPIPS noticeably decreases with more rays, interpreting that a higher number of warped rays ensures better perceptual quality. The performance for individual scenes in the synthetic dataset is shown in Tab. 7, showing that except for the “Cozyroom” scene, LPIPS decreases and overall performance peaks when $\mathcal{N}$ is 9. Analysis for the “Cozyroom” scene is conducted in Appendix D. As indicated in Tab. 8, for the real-world dataset, a larger $\mathcal{N}$ generally guarantees higher performance across most scenes. However, there are scenes where performance drops when $\mathcal{N}$ exceeds 9, suggesting that the optimal $\mathcal{N}$ might be 9. Even with $\mathcal{N}$ set to 5, which is the same condition to DP-NeRF [25] and Deblur-NeRF [31], SMURF outperforms them across all metrics. Furthermore, SMURF achieves higher performance with fewer warped rays than PDRF-10 [40], which set $\mathcal{N}$ at 10, validating the effectiveness of our proposed ideas.
Figure 6: Variation in performance with the number of warped rays.
Synthetic  $\mathcal{N}$ Factory CozyRoom Pool Tanabata Trolley
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
SMURF 4 25.13 0.7697 0.2127 32.46 0.9294 0.0407 33.04 0.8993 0.0825 28.38 0.9059 0.0825 28.57 0.9101 0.0890
SMURF 5 27.33 0.8381 0.2075 32.72 0.9315 0.0375 32.47 0.8904 0.0797 29.34 0.9222 0.0631 29.65 0.9241 0.0755
SMURF 6 26.74 0.8159 0.1897 31.74 0.9246 0.0381 32.56 0.8922 0.0778 29.22 0.9179 0.0604 28.99 0.9139 0.0599
SMURF 7 26.15 0.8010 0.1679 32.12 0.9269 0.0378 32.59 0.8926 0.0773 29.69 0.9269 0.0490 30.58 0.9348 0.0438
SMURF 8 29.87 0.8958 0.1057 32.48 0.9285 0.0379 32.34 0.8884 0.0779 29.91 0.9300 0.0436 30.30 0.9307 0.0397
SMURF 9 30.52 0.9065 0.0807 31.43 0.9198 0.0428 32.10 0.8849 0.0771 28.93 0.9184 0.0448 29.05 0.9136 0.0423
Table 7: Performance and training time of SMURF for individual scenes in the synthetic dataset, according to the number of warped rays.
Real-World  $\mathcal{N}$ Ball Basket Buick Coffee Decoration
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
SMURF 4 26.47 0.7424 0.1754 28.30 0.8653 0.0847 26.57 0.8270 0.1066 30.66 0.8676 0.1228 24.59 0.7972 0.1471
SMURF 5 26.63 0.7528 0.1549 27.96 0.8648 0.0741 26.03 0.8254 0.0947 30.93 0.8734 0.1172 24.22 0.7799 0.1497
SMURF 6 27.31 0.7678 0.1434 27.19 0.8463 0.0773 26.75 0.8315 0.0905 30.76 0.8731 0.1004 24.16 0.7744 0.1465
SMURF 7 27.68 0.7821 0.1325 29.24 0.8862 0.0617 27.05 0.8395 0.0867 30.67 0.8649 0.0931 24.97 0.8092 0.1220
SMURF 8 27.50 0.7760 0.1298 28.95 0.8842 0.0619 27.10 0.8409 0.0839 31.33 0.8879 0.0874 24.90 0.8114 0.1190
SMURF 9 27.16 0.7698 0.1315 28.52 0.8766 0.0631 26.92 0.8366 0.0813 31.41 0.8802 0.0870 24.12 0.7753 0.1405
Real-World  $\mathcal{N}$ Girl Heron Parterre Puppet Stair
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
SMURF 4 24.66 0.8410 0.1084 23.36 0.7144 0.1863 25.43 0.7644 0.1611 24.63 0.7496 0.1277 23.38 0.5406 0.1687
SMURF 5 24.73 0.8354 0.1028 23.47 0.7244 0.1717 25.22 0.7640 0.1474 24.69 0.7514 0.1239 25.22 0.6391 0.1171
SMURF 6 25.10 0.8473 0.0940 23.60 0.7386 0.1533 24.99 0.7569 0.1402 24.51 0.7493 0.1145 25.38 0.6371 0.0956
SMURF 7 25.43 0.8570 0.0884 23.66 0.7352 0.1457 25.22 0.7744 0.1302 24.70 0.7587 0.1076 25.42 0.6370 0.0846
SMURF 8 25.66 0.8592 0.0829 23.59 0.7317 0.1381 25.47 0.7825 0.1207 25.19 0.7702 0.1077 25.48 0.6421 0.0822
SMURF 9 25.56 0.8567 0.0822 23.81 0.7828 0.1333 24.91 0.7583 0.1216 24.71 0.7554 0.1035 25.31 0.6302 0.0777
Table 8: Performance and training time of SMURF for individual scenes in the real-world dataset, according to the number of warped rays.
Appendix C Additional Rendering Results
Additional rendering results are shown in Fig. 8 (for the synthetic dataset) and Figs. 9 and 10 (for the real-world dataset), demonstrating that our SMURF qualitatively offers the best quality when compared to the reference images.
Appendix D Analysis for Low-PSNR Scenes
Figure 7: Qualitative comparison of rendering results and error maps for the Cozyroom scene. DP-TensoRF is a model that applies the kernel proposed by DP-NeRF [25] to TensoRF [6].
In this section, we analyze individual scenes that showed slightly lower performance from the main paper . Notably, we present the error maps for the “Cozyroom” scene from the synthetic dataset for DP-NeRF [25], DP-TensoRF [25, 6], and SMURF in Fig. 7, where DP-TensoRF is a model that applies the blurring kernel proposed by DP-NeRF to TensoRF. The error map for DP-NeRF shows bright wall rendered cleanly without noise, whereas DP-TensoRF and SMURF exhibit noise on the wall. This indicates that the cause of noise is not the proposed CMBK but the inherent characteristics of the backbone model, TensoRF. As shown in Fig. 8, aside from the noise on the wall, the rendering results of SMURF show slightly sharper image quality in other areas except the wall. Moreover, the “Coffee”, “Parterre”, and “Puppet" scenes from real-world dataset, SMURF shows the best LPIPS score but somewhat lower PSNR. However, when comparing the rendering results in Figs. 9 and 10, it is observable that the results of SMURF are most similar to the reference images for these scenes.
Appendix E Limitations and Future Work
By adopting TensoRF [6], a 3D tensor factorization-based method, as our backbone, we ensure high quantitative performance, superior perceptual quality, and faster training. However, with the advent of 3D Gaussian Splatting [21], which allows for GPU-based rasterization instead of optimizing per ray, backbones that guarantee more faster training and rendering become available. Although our backbone may be slower compared to 3D Gaussian Splatting, applying the main idea of CMBK, continuous dynamics, to a rasterization-based method is expected to result in faster training and rendering. Furthermore, by demonstrating the applicability of continuous dynamics to the 3D scene deblurring, we anticipate the possibility of designing models that cover not only camera motion blur but also object motion blur, which is caused by the movement of objects within the scene.
Figure 8: Qualitative comparison for individual scenes in the synthetic dataset.
Figure 9: Qualitative comparison for individual scenes in the real-world dataset.
Figure 10: Qualitative comparison for individual scenes in the real-world dataset.

### v2

@@ 4.1 Implementation Details
Table 1: Quantitative results on synthetic and real-world scenes.
Methods Synthetic Scene Dataset Real-World Scene Dataset
Naive NeRF [15] 23.78 0.6807 0.3362 22.69 0.6347 0.3687
NeRF+MPR [23] 25.11 0.7476 0.2148 23.38 0.6655 0.3140
NeRF+PVD [20] 24.58 0.7190 0.2475 23.10 0.6389 0.3425
Deblur-NeRF [14] 28.77 0.8593 0.1400 25.63 0.7645 0.1820
PDRF-10* [17] 28.86 0.8795 0.1139 25.90 0.7734 0.1825
BAD-NeRF* [22] 27.32 0.8178 0.1127 22.82 0.6315 0.2887
DP-NeRF [12] 29.23 0.8674 0.1184 25.91 0.7751 0.1602
SMURF 30.98 0.9147 0.0609 26.52 0.7986 0.1013
@@
In this experiment, we utilize the camera motion blur dataset published by [14], which is divided into synthetic and real-world scenes. The synthetic scene dataset comprises five scenes synthesized using Blender [4], with multi-view cameras set up to render motion-blurred images. The real-world scene dataset consists of ten scenes captured with an actual camera. The blurry images are obtained by physically shaking the camera during the exposure time, and the camera poses are calibrated by COLMAP [19, 18].
@@ 4.2 Novel View Synthesis Results
We show the quantitative evaluation of our network, SMURF, in comparison to various baselines on the two datasets as shown in Tab. 1. We demonstrate superior quantitative performance across all the metrics. For qualitative results, we compare our results with previous 3D scene deblurring methods in Fig. 3.
Table 2: Ablations on embedding type and regularization strategies. “Emb.”, “O.S. Loss”, and “R.M.” refer to embedding type, the output suppression loss, and the residual momentum, respectively. “C.V” and “T” denote chrono-view and time embeddings.
Emb. O.S. R.M. Synthetic Dataset Real-World Dataset
PSNR SSIM LPIPS PSNR SSIM LPIPS
C.V 26.17 0.8186 0.0759 25.27 0.7576 0.1121
C.V 28.49 0.8727 0.0675 25.60 0.7748 0.1057
C.V 30.72 0.9052 0.0684 25.99 0.7846 0.1181
30.54 0.8995 0.0709 25.74 0.7755 0.1194
T. 30.94 0.9158 0.0610 26.34 0.7881 0.1102
C.V 30.98 0.9147 0.0609 26.52 0.7986 0.1013
@@ 4.3 Ablation Study
We conduct experiments by dividing the embedding types into three categories. The time embedding assumes that all images have the same exposure time without incorporating view information, while the chrono-view embedding applies view information, ensuring that all images have distinct exposure time. As shown in Tab. 2, the chrono-view embedding shows higher performance in real-world scenes, while its effect in synthetic scenes is minimal. This is attributed to the characteristics of the synthetic dataset created by Deblur-NeRF using Blender, where motion blur images are acquired by linearly interpolating between camera poses. In other words, all the images in the synthetic scenes set share the same exposure time. Our chrono-view embedding function assumes that all images have varying exposure times, leading to its minimal effect on the synthetic dataset. However, the real-world dataset consists of images captured with an actual camera, where the speed of camera motion is not constant during the exposure time. Therefore, applying this function to the real-world scenes yield improved performance due to these variations.
@@
We conduct various experiments for regularization strategies as shown in Tab. 2. The output suppression loss yields higher performance as the changes in rays are prevented from diverging since the estimated poses also share the same derivative function $f$. For residual momentum, the sequential camera pose does not deviate significantly from the previous one. Tab. 2 shows that applying residual momentum shows better performance in terms of LPIPS.
@@ 5 Conclusion
We have proposed SMURF, a novel approach that sequentially models camera motion for reconstructing sharp 3D scenes from motion-blurred images. SMURF incorporates a kernel for estimating sequential camera motions, named the CMBK with by two regularization techniques: residual momentum and output suppression loss. We model the 3D scene using TensoRF, which allows for the integration of blurred information and sharp information within adjacent voxels. SMURF significantly outperform previous works quantitatively and qualitatively.